Convert the following masses into their corresponding number of moles.




6. 23.5 g of sodium chloride




7. 0.778 g of sodium cyanide




8. 0.250 g of water




9. 169.45 g of calcium acetate




10. 79.9 g of potassium permanganate

1
Expert's answer
2022-03-12T02:01:11-0500

6. "\frac{23.5}{58.44}=0.4021" moles


7. "\frac{0.778}{49.01}=0.0159" moles


8. "\frac{0.250}{18.02}=0.0139" moles


9. "\frac{169.45}{158.17}=1.0713" moles


10. "\frac{79.9}{158.03}=0.5056" moles